OpenStudy (anonymous):

h(x)=x^2+10x+16 can anyone complete the square and explain the step?

ganeshie8 (ganeshie8):

whats the coefficient of x ?

OpenStudy (anonymous):

10 right?

ganeshie8 (ganeshie8):

yes, take half of it : 10/2 = 5

ganeshie8 (ganeshie8):

square it, then add and subtract

ganeshie8 (ganeshie8):

h(x) = x^2 + 10x + 16 = x^2 + 10x + 5^2 - 5^2 + 16

ganeshie8 (ganeshie8):

recall the identity : `a^2 + 2ab + b^2 = (a+b)^2`

ganeshie8 (ganeshie8):

h(x) = x^2 + 10x + 16 = x^2 + 10x + 5^2 - 5^2 + 16 = (x+5)^2 - 5^2 + 16

ganeshie8 (ganeshie8):

we're done ! just simplify the last two terms ^

OpenStudy (anonymous):

So (x+5)^2 -9 is the answer? I have to find 2 complex solutions, is this one?

ganeshie8 (ganeshie8):

thats `completing the square` part of the problem

ganeshie8 (ganeshie8):

to find the solutions, set it equal to 0 and solve x

OpenStudy (anonymous):

From the complete the square part of from the beginning?

ganeshie8 (ganeshie8):

(x+5)^2 -9 = 0 (x+5)^2 = 9 x+5 = +-3

ganeshie8 (ganeshie8):

solve x^

OpenStudy (anonymous):

So its just 2/-2?

ganeshie8 (ganeshie8):

x+5 = +-3 x+5 = 3 ; x+5 = -3 x = 3-5 ; x = -3-5 x = -2 ; x = -8

OpenStudy (anonymous):

Ohhh, right, oops. But those arent complex solutions are they..? What kid of equation will give us one?

ganeshie8 (ganeshie8):

to get complex solutions, you need an equation with below condition : b^2 - 4ac < 0

ganeshie8 (ganeshie8):

try below : h(x) = x^2 + 2x + 2

OpenStudy (anonymous):

So we'd have x^2+2x+SQRT2^2-SQRT2^2+2?

ganeshie8 (ganeshie8):

here are the steps to complete the square : 1) take half of x coefficient 2) square it, add and subtract 3) complete the square using the identity `a^2 + 2ab + b^2 = (a+b)^2` 4) simplify the tail

ganeshie8 (ganeshie8):

h(x) = x^2 + 2x + 2 1) take half of x coefficient what do you get ?

OpenStudy (anonymous):

1

ganeshie8 (ganeshie8):

2) square it, add and subtract what do you get ?

OpenStudy (anonymous):

x^2+2x+1^2-1^2

OpenStudy (anonymous):

+2

ganeshie8 (ganeshie8):

don't forget the constant term 2 in the end ^

ganeshie8 (ganeshie8):

yes ! h(x) = x^2 + 2x + 2 = x^2 + 2x + 1^2 - 1^2 + 2

OpenStudy (anonymous):

And then we have 1+4+4=(a+b)^2 When we do that A^2+2ab+b^2=(a+b)^2

ganeshie8 (ganeshie8):

yes, a = x, b = 1

ganeshie8 (ganeshie8):

h(x) = x^2 + 2x + 2 = x^2 + 2x + 1^2 - 1^2 + 2 = (x+1)^2 -1^2 + 2

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

So really, it is just (x+1)^2+1?

ganeshie8 (ganeshie8):

Yep ! set it equal to 0 and solve x

OpenStudy (anonymous):

So what do I do after -1=(x+1)^2?

ganeshie8 (ganeshie8):

take sqrt both sides

OpenStudy (anonymous):

And then 1i=x+1 goes to x=1i-1?

ganeshie8 (ganeshie8):

very good, but looks u made a small mistake. \(\large (x+1)^2 = -1 \) take sqrt both sides : \(\large (x+1) = \pm \sqrt{-1} \) \(\large x+1 = \pm i \) \(\large x = ?\)

OpenStudy (anonymous):

x=1+/- i?

OpenStudy (anonymous):

-1*

ganeshie8 (ganeshie8):

Yes !

ganeshie8 (ganeshie8):

\(\large x = -1 \pm i\)

OpenStudy (anonymous):

Thank you for the walkthrough, I really appreciate it!

ganeshie8 (ganeshie8):

np, you're welcome :)